OpenStudy (anonymous):

A right circular cone has a volume of 192pi cm cubed. Find the radius and the height that minimize its surface area.

OpenStudy (solomonzelman):

\[V=\pi r ^2 \frac{h}{3}\]

OpenStudy (solomonzelman):

\[192=\pi r ^2 \frac{h}{3}\]

OpenStudy (solomonzelman):

\[576=\pi h r ^2\]

OpenStudy (solomonzelman):

\[h=\frac{576}{\pi}r^{-2}\]

OpenStudy (solomonzelman):

I am thinking of this, then direivative for critical numbers,